# End-to-End Learning of Transmitter and Receiver Filters in Bandwidth Limited Fiber Optic Communication Systems

### v1

@@ I Introduction
Current solutions, for mitigating ISI are predominantly focused on receiver-side equalization, i.e. the process of removing the distortion introduced by the system, which can be done using an adaptive linear filter, denoted a feed-forward equaliser (FFE) [4]. In systems with non-linear distortion due to hardware imperfections and fiber dispersion, non-linear compensation can be employed. Examples include non-linear adaptive filters such as the Volterra equalizer [5] and neural networks (see [6] for an overview). Some effort has also gone into mitigating distortion from the transmitter side, either through pulse-shaping [7], i.e. designing spectrally efficient filters to minimize ISI, or predistortion [8, 9] and constellation shaping [10] to linearize the transmitter response. Typically these compensation mechanisms are designed or optimized independently of each other, which may lead to suboptimal performance.
End-to-end learning (E2E)[11, 12] has emerged as a framework for jointly optimizing all adaptable DSP blocks in the system on both the transmitter and receiver side. In [12] this was shown for an IM/DD system, where both transmitter and receiver was modeled by a feed-forward neural network using an autoencoder-style structure. The system was trained offline (through numerical simulation) using the backpropagation algorithm given a known pilot sequence, and evaluated experimentally with a fine-tuning step for the receiver. This enabled information rates above 42 Gb/s beyod 40 km of transmission distance without amplification. E2E has since been used in the context of geometric constellation shaping [13], wavelength division multiplexing systems [14], learning parameters in directly modulated laser transmission [15] and learning the pulse shape and receiver demapper in a 6G wireless application [16], to mention a few examples.
@@
In this paper we focus on the setting where the pulse-shaper and receiver filter are implemented as FIR filters, which is of great importance due to low complexity of implementation. We numerically investigate the use of E2E for jointly optimizing the pulse-shaper and receiver filter to combat ISI due to bandwidth limitations of the digital-to-analogue (DAC) and analogue-to-digital (ADC) converter, as well as chromatic dispersion, in a 100 GBd four pulse amplitude modulation (4-PAM) communication link. We first show the benefits of joint pulse-shaper and receiver filter optimization in a simple additive white Gaussian noise (AWGN) system with bandwidth limitation, where zero ISI can be reached for fairly short filter lengths. Next, we show the E2E framework’s benefits in a wavelength division multiplexing (WDM) scheme, where the learned filters must adhere to the bandwidth restrictions imposed by the presence of interfering channels. Finally, we show the benefits of joint optimization for a 1 km WDM IM/DD link with an electro-absorption modulator (EAM) that exhibits some degree of nonlinearity.
@@ II-B Additive White Gaussian Noise Channel with Bandwidth Limitation
During training (see Figure 1a), the output of the receiver filter, $\hat{\bm{x}}$, is compared to the true symbol sequence $\bm{x}$ using MSE, as described in the previous section. The $\hat{\bm{x}}$ and $\bm{x}$ are coarsely aligned before caculating the cost function, by a priori calculating the average group-delay of the two Bessel filters in the passband and compensating for that in the MSE calculation.
@@ II-C Multilevel Intensity Modulated and Direct Detected System
$E_{\text{ideal}}(t)=\sqrt{P_{\text{in}}\cdot V_{\text{m}}(t)},$ (2)
@@
$A_{\text{EAM}}(t)=\sqrt{P_{\text{in}}\cdot 10^{-\alpha\left(V_{\text{m}}(t)\right)/10}},$ (3)
@@
in which $\alpha$ is the linewidth enhancement factor that controls the amount of chirp. The voltage-to-optical amplitude relationship and the absorption curve for the EAM used in the simulations can be seen in Figure 3. Chromatic dispersion and fiber attenuation, associated with the SMF, is modeled using a standard frequency domain filter [22]. We use the laser and fiber parameters from [10] (cf. Table I). The signal is then detected using a square-law detection, with thermal noise and shot noise both modeled as white Gaussian random variables with variances $\sigma_{\text{thermal}}^{2}$ and $\sigma_{\text{shot}}^{2}$, respectively. The noise variances are parameterized as,
@@
Modulator Laser power ($P_{\text{in}}$) -
@@ III-A Additive White Gaussian Noise Channel with Bandwidth Limitation
We compare the four learning methods on the AWGN channel described in section II-B. We vary the SNR and plot the SER for different filter lengths, which can be seen in Figure 5. Note that all filter optimization is carried out at an SNR of 12 dB. The theoretical curve is based on an AWGN channel without bandwidth limitation. The joint optimization framework (PS & RF) comes extremely close to the theoretical limit even for very short filter lengths. The penalty to the theory almost vanishes with a filter length of $N=25$. All other methods incur substantial penalties irrespective of the filter length. The baseline (RRC & FFE) has a slight advantage over the other single-sided methods, which is attributed to the extra taps in the FFE. However, this advantage reduces as we increase the filter lengths.
@@
in which $T$ is the sampling interval. In a ISI-free system the quantity $B(f)$ is constant over all frequencies [4]. We see in Figure 6b that the joint optimization scheme achieves the flattest curve, which explains the near-optimal symbol-error performance in the AWGN channel (see Figure 5). Furthermore, we see in Figure 6a that the methods that learn the pulse-shaper (PS and PS & RxF) utilize a much wider range of frequencies compared to the RRC filter. This is due to the unconstrained nature of our cost function in the optimization problem [14].
@@ III-C Intensity Modulated and Direct Detected System
Figure 10: IM/DD system with EAM and channel spacing 200 GHz: SER vs $V_{\text{pp}}$ in DAC for different fiber and filter lengths.
Figure 11: IM/DD system with EAM modulator: SER vs WDM channel spacing for different fiber and filter lengths with $V_{\text{pp}}=4.0$ in DAC.
@@
Finally, we investigate the learning framework with a non-linear modulator, namely the EAM (see input-output characteristic in Figure 3b). In all simulations with the EAM, in addition to the filters, we optimize $g_{\text{dac}}$ and the bias, $V_{\text{b}}$, in the DAC such that each method has the flexibility to utilize the modulator range in the best way possible. We initialized the bias in modulator to $V_{\text{b}}=-1.0$ for all simulations as we found that to yield consistently good results. The results of the numerical simulations can be seen in Figure 10 and 11. In the B2B condition in Figure 10, due to the non-linear characteristic of the EAM, we see that the SER curve flattens as we increase the peak-to-peak voltage. The relative performance between the methods is consistent with previous results, namely that the joint-learning framework outperforms the ”single-side” frameworks. When we increase the fiber length to 1 km, we see a dramatic drop in performance and more linear behaviour as a function of $V_{\text{pp}}$. We primarily attribute this to the phase noise introduced by the chirp in the modulator that significantly impacts the performance of all methods. Similar to the other simulations, the channel spacing in the WDM needs to be at least 125 GHz for the end-to-end learning framework to have an advantage over the other methods (see Figure 11). However, we note that the for a fiber length of 1 km all methods struggle to reach good SER performance and the difference over channel spacing is relatively low (especially for short filters). An example of eyediagrams after receiver filtering in the 1 km case can be seen in Figure 12. The joint optimization (PS & RxF) achieves the clearest eye-opening compared to the other methods.
@@ IV Conclusion
We investigated the use of end-to-end learning for joint optimization of pulse-shaping and reciver-side FIR filters to mitigate ISI in bandwidth-limited communication channels. By implementing the channels in PyTorch, with the use of automatic differentiation we numerically investigated the performance gains in two types of systems; an AWGN channel with bandwidth-limitation and an IM/DD short-reach link with a non-linear modulator. We compared joint optimization to their respective single-sided counterpart, fixing one side to a RRC filter and learning the other side. We consistently found that joint optimization outperformed the other methods across both channels and SNRs. In the IM/DD system, we found support for short filter lengths ($N_{\text{taps}}=15$) being able to respect the KP4 forward-error correction threshold, which none of the other methods came close. This points toward joint-optimization methods being able to deliver considerable savings in terms of operating IM/DD links, at the (minor) cost of some calibration and optimization.

### v2

@@ I Introduction
Current solutions, for mitigating ISI are predominantly focused on receiver-side equalization, i.e. the process of removing the distortion introduced by the system, which can be done using an adaptive linear filter, denoted a feed-forward equaliser (FFE) [4]. Feedback filter structures are often used in conjunction with the FFE, either after the FFE at the receiver, denoted decision feedback equalizer (DFE) [5], or at the transmitter, denoted Tomlinson-Harashima Precoding (THP) [6]. In systems with non-linear distortion due to hardware imperfections and fiber dispersion, non-linear compensation can be employed. Examples include non-linear adaptive filters such as the Volterra equalizer [7] and neural networks (see [8] for an overview). Some effort has also gone into mitigating distortion from the transmitter side, either through pulse-shaping [9], i.e. designing spectrally efficient filters to minimize ISI, or predistortion [10, 11] and constellation shaping [12] to linearize the transmitter response. Typically these compensation mechanisms are designed or optimized independently of each other, which may lead to suboptimal performance.
End-to-end learning (E2E)[13, 14] has emerged as a framework for jointly optimizing all adaptable DSP blocks in the system on both the transmitter and receiver side. In [14] this was shown for an IM/DD system, where both transmitter and receiver were modeled by a feed-forward neural network using an autoencoder-style structure. The system was trained offline (through numerical simulation) using the backpropagation algorithm given a known pilot sequence, and evaluated experimentally with a fine-tuning step for the receiver. This enabled information rates above 42 Gb/s beyond 40 km of transmission distance without amplification. E2E has since been used in the context of geometric constellation shaping [15], joint geometric and probabilistic constellation shaping [16], wavelength division multiplexing systems [17], learning parameters in directly modulated laser transmission [18] and learning the pulse shape and receiver demapper in a 6G wireless application [19], to mention a few examples.
@@
In this paper we focus on the setting where the pulse-shaper and receiver filter are implemented as FIR filters, which is of great importance due to low complexity of implementation. We numerically investigate the use of E2E for jointly optimizing the pulse-shaper and receiver filter to combat ISI due to bandwidth limitations of the digital-to-analogue (DAC) and analogue-to-digital (ADC) converter, as well as chromatic dispersion, in a 100 GBd four pulse amplitude modulation (4-PAM) communication link. We first show the benefits of joint pulse-shaper and receiver filter optimization in a simple additive white Gaussian noise (AWGN) system with bandwidth limitation, where zero ISI can be reached for fairly short filter lengths. Next, we show the E2E framework’s benefits in a wavelength division multiplexing (WDM) scheme, where the learned filters must adhere to the bandwidth restrictions imposed by the presence of interfering channels. We then show the benefits of joint optimization for a 1 km WDM IM/DD link with an electro-absorption modulator (EAM) that exhibits some degree of nonlinearity. In the non-linear case, we additionally compare our proposed joint optimization to a flexible Volterra equalizer with a large number of parameters to investigate the theoretical limits of this type of optimization. Finally, we evaluate the robustness of the learned filters to fiber non-linearities using the split-step Fourier method (SSFM).
@@ II-B Additive White Gaussian Noise Channel with Bandwidth Limitation
During training (see Figure 1a), the output of the receiver filter, $\hat{\bm{x}}$, is compared to the true symbol sequence $\bm{x}$ using MSE, as described in the previous section. The $\hat{\bm{x}}$ and $\bm{x}$ are coarsely aligned before calculating the cost function, by a priori calculating the average group-delay of the two Bessel filters in the passband and compensating for that in the MSE calculation.
@@ II-C Multilevel Intensity Modulated and Direct Detected System
$E_{\text{ideal}}(t)=\sqrt{P_{\text{in}}^{(\text{ideal})}\cdot V_{\text{m}}(t)},$ (2)
@@
$A_{\text{EAM}}(t)=\sqrt{P_{\text{in}}^{(\text{EAM})}\cdot 10^{-\alpha\left(V_{\text{m}}(t)\right)/10}},$ (3)
@@
in which $\alpha$ is the linewidth enhancement factor that controls the amount of chirp. The voltage-to-optical amplitude relationship and the absorption curve for the EAM used in the simulations can be seen in Figure 3. For both types of modulators, we keep the laser power fixed at a given level (cf. Table I) and only change $V_{\text{pp}}$ in our simulations. Chromatic dispersion and fiber attenuation, associated with the SMF, is modeled using a standard frequency domain filter [25]. We use the laser and fiber parameters from [12] (cf. Table I). The signal is then detected using a square-law detection, with thermal noise and shot noise both modeled as white Gaussian random variables with variances $\sigma_{\text{thermal}}^{2}$ and $\sigma_{\text{shot}}^{2}$, respectively. The noise variances are parameterized as,
@@
Modulator Laser power ($P_{\text{in}}^{(\text{ideal})}$) -13.0 dBm
Laser power ($P_{\text{in}}^{(\text{EAM})}$) -6.0 dBm
@@ III-A Additive White Gaussian Noise Channel with Bandwidth Limitation
We compare the four learning methods on the AWGN channel described in section II-B. We vary the SNR and plot the SER for different filter lengths, which can be seen in Figure 5. Note that all filter optimization is carried out at an SNR of 12 dB. The choice of SNR does not influence the learned filters as the channel transfer function is linear. The theoretical curve is based on an AWGN channel without bandwidth limitation. The joint optimization framework (PS & RF) comes extremely close to the theoretical limit even for very short filter lengths. The penalty to the theory almost vanishes with a filter length of $N=25$. All other methods incur substantial penalties irrespective of the filter length. The baseline (RRC & FFE) has a slight advantage over the other single-sided methods, which is attributed to the extra taps in the FFE. However, this advantage reduces as we increase the filter lengths.
@@
in which $T$ is the sampling interval. In an ISI-free system the quantity $B(f)$ is constant over all frequencies [4]. We see in Figure 6b that the joint optimization scheme achieves the flattest curve, which explains the near-optimal symbol-error performance in the AWGN channel (see Figure 5). Furthermore, we see in Figure 6a that the methods that learn the pulse-shaper (PS and PS & RxF) utilize a much wider range of frequencies compared to the RRC filter. This is due to the unconstrained nature of our cost function in the optimization problem [17].
@@ III-C Intensity Modulated and Direct Detected System
Figure 10: IM/DD system with EAM and channel spacing 200 GHz: SER vs $V_{\text{pp}}$ in DAC for different fiber and filter lengths. Note, the Volterra* methods has many more free parameters than the other methods and are only meant as a theoretical comparisons. The shaded area represents a 95% confidence interval estimated over 5 random restarts of the simulation.
Figure 11: IM/DD system with EAM modulator: SER vs WDM channel spacing for different fiber and filter lengths with $V_{\text{pp}}=4.0$ in DAC. Note, the Volterra* methods has many more free parameters than the other methods and are only meant as a theoretical comparisons. The shaded area represents a 95% confidence interval estimated over 5 random restarts of the simulation.
@@
Finally, we investigate the learning framework with a non-linear modulator, namely the EAM (see input-output characteristic in Figure 3b). In all simulations with the EAM, in addition to the filters, we optimize $g_{\text{dac}}$ and the bias, $V_{\text{b}}$, in the DAC such that each method has the flexibility to utilize the modulator range in the best way possible. We initialized the bias in modulator to $V_{\text{b}}=-1.0$ for all simulations as we found that to yield consistently good results. Due to the non-linear modulator characteristic, we additionally compare the filter optimization methods with a second order Volterra equalizer. We set the number of taps in the first order filter to 101 and 45 taps in the second order, yielding 1136 free parameters. We use the non-linear equalizer in two configurations — 1) together with a fixed RRC filter on the Tx side and a fixed RRC filter on the Rx side denoted RRC & Volterra* and 2) together with a learnable pulse-shaper and fixed RRC filter on the Rx side denoted PS & Volterra*. The two Volterra* methods can viewed as a best-case comparison, where we have not taken into account computational complexity constraints. The results of the numerical simulations can be seen in Figure 10 and 11. In the B2B condition in Figure 10, due to the non-linear characteristic of the EAM, we see that the SER curve flattens as we increase the peak-to-peak voltage, except for the Volterra* methods that continue to improve. The relative performance between the methods is consistent with previous results, namely that the joint-learning framework outperforms the ”single-side” frameworks. One might expect that the Volterra equalizer alone (RRC & Volterra*) should be enough to reach the optimum performance, with a clear advantage in the number of free parameters over the other standard filtering methods (PS, RxF and PS & RxF). We attribute this to the ISI of the system stemming from bandwidth limitation being the limiting factor, and not the non-linearity of the system. In a channel where the non-linearities are more dominant we would expect the Volterra alone to yield better results, but this was not the focus of the paper. If we additionally optimize the pulse-shaper, as represented by the PS & Volterra method, we observe further gains in terms of performance compared to the linear approach (PS & RxF). When we increase the fiber length to 1 km, we see a dramatic drop in performance and more linear behavior as a function of $V_{\text{pp}}$. We primarily attribute this to the phase noise introduced by the chirp in the modulator that significantly impacts the performance of all methods. Similar to the other simulations, the channel spacing in the WDM needs to be at least 125 GHz for the end-to-end learning framework to have an advantage over the other methods (see Figure 11). However, we note that the for a fiber length of 1 km all methods struggle to reach good SER performance and the difference over channel spacing is relatively low (especially for short filters). As in the simulations with the ideal modulator, the ”receiver-side only” methods (RxF, RRC & FFE and RRC & Volterra*) all incur a penalty when increasing the filter length from 25 to 65. As already discussed in the previous simulation, we attribute this to the clipping constraint in the DAC and the high PAPR of the RRC filter at the Tx. An example of eyediagrams after receiver filtering in the 1 km case can be seen in Figure 12. The joint optimization (PS & RxF) achieves the clearest eye-opening compared to the other methods.
@@
III-C1 Filter Robustness Evaluation with Non-linear Fiber Model
Robustness of end-to-end frameworks will always be of interest as deviations from the simulation model will lead to performance degradation, unless specifically accounted for during training [14, 15]. To investigate the learned pulse-shaper and receiver filters’ robustness and generalizability, we devised a small simulation in which the channel conditions are altered during evaluation.
The E2E systems under test, are trained on the IM/DD WDM system described above with a 1 km SMF with parameters described in Table I. We used 3 WDM channels with a channel spacing of 150 GHz. During evaluation, we introduce a non-linear fiber with the split-step Fourier method (SSFM) with a non-linearity coefficient of $\gamma=1.3$ [1/W/km]. The step size was set to $0.25$ km and we did not use any amplification. We introduce 2 additional interferer channels, bringing the total number of channels up to 5, and sweep the launch power that enters the fiber. The results of this simulation for PS, RxF and PS & RxF can be seen in Figure 13. Overall, all the methods follow the same trend, namely that the SER decreases until the non-linear distortion from the fiber becomes dominant and we start to see the SER increase. The joint optimization method is across the board the best method, even in the highly non-linear regime, however, it also sees the biggest increase in SER compared relative to the optimum performance.
Figure 13: Robustness simulation: The plots show the evaluation performance of the IM/DD WDM (5 channel) systems with a non-linear fiber ($\gamma=1.3$ [1/W/km]) as a function of laser input power ($P_{\text{in}}^{(\text{EAM})}$). The systems were trained with 3 channels on a standard SMF at $P_{\text{in}}^{(\text{EAM})}=-6.0$ dBm, which in launch power is indicated by the dashed black vertical line. During both training and evaluation the channel spacing was set to 150 GHz.
@@ IV Conclusion
We investigated the use of end-to-end learning for joint optimization of pulse-shaping and reciver-side FIR filters to mitigate ISI in bandwidth-limited communication channels. By implementing the channels in PyTorch, with the use of automatic differentiation we numerically investigated the performance gains in two types of systems; an AWGN channel with bandwidth-limitation and an IM/DD short-reach link with a non-linear modulator. We compared joint optimization to their respective single-sided counterpart, fixing one side to a RRC filter and learning the other side. We consistently found that joint optimization outperformed the other methods across both channels and SNRs. In the IM/DD system, we found support for short filter lengths ($N_{\text{taps}}=15$) being able to respect the KP4 forward-error correction threshold which none of the other linear methods came close to. We hypothesize that joint optimizations’ superiority comes from the additional degrees of freedom at the Tx, which results in the channel response, including hardware components, being more efficiently invertible. This points toward joint-optimization methods being able to deliver considerable savings in terms of operating IM/DD links, at the (minor) cost of some calibration and optimization.